OpenStudy (anonymous):

a right angle triangle has one length 5m and one 10m what is the length of the 3rd side?

OpenStudy (anonymous):

do you know pythagoras rule? a² + b² = c² (the letters are the length of the sides). Can you solve it now?

OpenStudy (anonymous):

Do you know if the 10 is the longest side or are you solving for the longest side?

OpenStudy (anonymous):

solving the longest side

OpenStudy (anonymous):

Yes he is right use Pythagoras theorem here to find the third side. p^2+b^2=h^2 Where p is perpendicular b is base h is height

OpenStudy (anonymous):

Using Simon's help above and a=5 and b=10 we should get: 5^2+10^2=c^2 25+100=c^2 125=c^2 c = sqrt(125) or about 11.18